OpenStudy (anonymous):

the equation y''+k^2y=0 is known as a different equation, since it involves a function y and its derivatives. this equation is widely used in applications to model physical systems that oscillate. for example, y could represent the voltage in an electrical circuit, or the position of a vibrating airplane wing. show that the function y=Asin(kx+B) satisfies the differential equation for all values of the constants A,B, AND k.

OpenStudy (blockcolder):

y'=Ak cos(kx+B) y''=-A(k^2) sin(kx+B) y''+k^2y=-Ak^2 sin(kx+B)+Ak^2sin(kx+B)=0

OpenStudy (anonymous):

thanks alot blockcolder

OpenStudy (blockcolder):

No problemo. :D